A U.S. census bureau pollster noted that in 349 random households surveyed, 220 occupants owned their own home. What is the 95% confidence interval estimate of the proportion of American households who own their own home?

Question 9 options:

1)

( 0.31898 , 0.42027 )

2)

( 0.60453 , 0.65621 )

3)

( 0.57973 , 0.68102 )

4)

( 0.58787 , 0.67287 )

5)

( -0.57973 , 0.68102 )
Answer #1

Solution: Given that Sample Size 0 - 349 Sample proportion X = 220 So, population poopootion 220 349 - 01630 3125 1-ệ : 03662ME=1196 * 10.00066763 = 1.96*0.0258385 20:050643 954. Confidence inteoval is Ê IME 70.6303725 + 0-050643. 7(0.5797295; 0.6810